OpenStudy (bossimbacon):

How many solutions does the equation have? 2(2x+5)=4(x+3) A. one solution B. infinite solutions C. no solution @cj49

OpenStudy (bossimbacon):

@Haseeb96

OpenStudy (cwrw238):

2(2x+5)=4(x+3) 4x + 20 = 4x + 12 subtract 4x from both sides 20 = 12 No solution

OpenStudy (haseeb96):

It has no solution

OpenStudy (cj49):

^

OpenStudy (bossimbacon):

thanks can i ask doe more?

OpenStudy (cwrw238):

* should be 4x + 10 10 = 12 still no solution

OpenStudy (bossimbacon):

How many solutions does the equation have? a+5=1/5(5a+25) A. one solution B. infinite solutions C. no solution

OpenStudy (cwrw238):

a+5=1/5(5a+25) a + 5 = a + 5 you can assign any value you like to a so infinite solutions

OpenStudy (bossimbacon):

thanks^ one few more?

OpenStudy (bossimbacon):

How many solutions does the equation have? 3(4x-3)-7x=5x-9 A. one solution B. infinite solutions C. no solution

OpenStudy (cwrw238):

can you try this yourself? first distribute the 3 over the parentheses; 12x - 9 - 7x = 5x - 9 can you continue?

OpenStudy (cwrw238):

work out 12x - 7x = ?